# Learned frequency-domain scattered wavefield solutions using neural operators

### v2

@@ III-D Generalization test with reference frequency:
As shown in [28], FNO faces aliasing problems across different resolution scales, yielding large errors when we apply the pre-trained neural operator to predict wavefields for an unseen resolution, like when we increase the simulation domain. So, we extend our testing to evaluate the proposed method’s generalization capabilities on an unseen sample for a larger model, utilizing the concept of reference frequency. The neural network was initially trained on velocities covering an area of 6.4$\times$6.4 km2 with a resolution of 256$\times$256. We then conducted a generalization test on a velocity model spanning 12.8$\times$12.8 km2. with the same spatial interval. Due to the inherent resolution independence of the FNO, it seamlessly outputs the wavefield for the larger domain, which we refer to as ’direct prediction’, as depicted in Figure 9. Although the main components of the wavefield have been predicted by the FNO, their relative errors are still large. We calculate the relative L2 norm, which is the Euclidean distance from the prediction to the reference solutions divided by the Euclidean norm of the reference solutions, and the relative errors for direct predictions are 0.520 for the real part and 0.518 for the imaginary part of the wavefields. It is hard to obtain a high-accuracy prediction for a large unseen domain directly. Crucially, leveraging the reference frequency concept [18] allows us to upscale the 4.78 Hz wavefield in the large domain (12.8$\times$12.8 km2) to an equivalent 9.56 Hz wavefield in the smaller domain (6.4$\times$6.4 km2). This provides us with physical support to predict high-frequency wavefields for a large-scale velocity model rather than a direct prediction of high-resolution wavefields. As illustrated in Figure 10, this scaled wavefield demonstrates enhanced accuracy compared to direct predictions. The relative errors calculated at the resolution of 256$\times$256 between the predictions of our method without interpolation and the reference results, shown in Figure 9, are 0.052 for the real part and 0.054 for the imaginary part of the wavefield. After bilinear interpolation of the prediction using our method, the errors at a resolution of 512$\times$512 are 0.096 for the real part and 0.097 for the imaginary part of the wavefield. Though the errors slightly increase, as shown in Figure 10, the predictions using our method are still close to the reference results.
@@
[…]
IV Discussions
@@
The proposed method has shown good performance on the OpenFWI and realistic velocity models. It frees us at inference time from numerical simulations, in which the computational complexity is high. Taking the test on realistic models as an example, the numerical simulation for each velocity and frequency, executed in parallel on an Intel(R) Xeon(R) Gold 6230R CPU @2.10GHz, at a resolution of 139$\times$139 requires 1.27 seconds, and that at a resolution of 256$\times$256 and 512$\times$512 require 2.79 and 8.64 seconds, respectively. Using our method, the elapsed time on an NVIDIA A100 GPU (we do not fully utilize the GPU due to the relatively small input size and batch size) is 0.008 seconds for the resolution of 139$\times$139, 0.012 seconds for the resolution of 256$\times$256, and 0.031 seconds for the resolution of 512$\times$512. The main computationally intensive part of the proposed method is the training. In our experiments, it took 15.75 hours for training the neural operator for the ”CurveVelA velocity models, and 29.25 hours for the realisitc velocity models. Noteworthy, all of the training are performed offline. When trying to apply our method, indeed the neural operator-based wavefield solutions, on the downstream tasks such as reverse time migration or full waveform inversion, which requires massive simulations on different frequencies and source locations, the speed-up ratio of our method compared to conventional numerical methods will significantly increase, which will easily justify the expensive training. Taking multiple frequencies simulations on a resolution of 139$\times$139 as an example, the elapsed times for learned wavefield solutions and numerical finite-difference method with respect to the number of frequencies are shown in Figure 13. Since the learned wavefield solutions make it easy to parallel the prediction without additional modification (like parallel computing required for the numerical method to handle multiple frequencies simultaneously), instead, we just increase the batch size of the input tensors. It is obvious that within 16 simulations, the elapsed time for the learned wavefield solutions did not exceed 1 s, while the elapsed time for the numerical simulation without additional modification increased linearly. Due to the saturation of computational resources and memory bandwidth, the cost of neural operator-based wavefield solutions is slightly increased, causing a non-linear increase in the speed-up ratio. In summary, at inference, it is much faster than the conventional numerical solvers. Those speedups come from two main aspects: (1) we avoid repeated simulations; (2) we can easily utilize the latest hardware and software advances of accelerated computing [29] because we build the algorithm on machine learning. However, we recognize the challenges associated with the preparation of data-label pairs required for training the neural operator, particularly as the resolution of training samples increases.
@@
Figure 11: The relative L2 norm errors for each class of models. Regarding the sample index from 0 to 120, each of the 24 consecutive numbers represents a velocity model, and each of the three consecutive numbers represents a frequency, which increases in sequence. Each frequency value corresponds to three different shots.
@@
Figure 13: The efficiency comparison in terms of elapsed time for the learned wavefield solutions and the numerical finite-difference method on a resolution of 139$\times$139 with respect to different number of frequencies.
@@
Finally, to further highlight the effectiveness of the proposed method on different neural network architectures, we compare the different approaches such as such as conventional inputs with full wavefield outputs (marked as ”Conventional in full out”), conventional inputs with scattered wavefield outputs (marked as ”Conventional in residual out”), our method with full wavefield outputs, and our method with scattered wavefield outputs on different neural network architectures. We pick ResNet [22] and Unet [23] for the study. The results are shown in Tables II and III. The results demonstrate that our method is generally effective with other neural networks.
TABLE II: The training cost and accuracy comparison between different approaches, using ResNet as backbone
@@
TABLE III: The training cost and accuracy comparison between different approaches, using Unet as backbone
@@
V Conclusions
We presented a novel implementation for predicting the wavefield using the Fourier Neural Operator (FNO) by incorporating the analytically derived background wavefield as input. As a result, the source location and frequency are embedded in the background wavefield, facilitating compact and effective feature engineering. This methodology enables efficient surrogate modeling across various velocities, frequencies, and source locations. Our tests on both the OpenFWI dataset and some realistic velocity models have demonstrated the high accuracy and effectiveness of the proposed method. Additionally, incorporating a reference frequency strategy significantly enhances the accuracy of our model in generalizing to larger, unseen domain, marking a substantial advancement in the field of frequency-domain wavefield modeling.
VI Acknowledgements

### v3

@@ 3.4 Generalization test with reference frequency
While FNO supports zero-shot super-resolution simulation (Li et al., 2020), as shown in Raonić et al. (2023), it faces aliasing problems across different resolution scales, yielding large errors when we apply the pre-trained neural operator to predict wavefields for an unseen resolution. Beyond super-resolution, where physical simulations are performed on the same spatial domain but with finer intervals, generalizing to larger domains is challenging. In this scenario, the spatial resolution remains consistent with the training data, but the overall coverage area expands. So, we extend our testing to evaluate the proposed method’s generalization capabilities on an unseen sample for a larger model, utilizing the concept of reference frequency. The neural network was initially trained on velocities covering an area of 6.4$\times$6.4 km2 with a resolution of 256$\times$256. We then conducted a generalization test on a velocity model spanning 12.8$\times$12.8 km2, with the same spatial interval. Due to the inherent resolution independence of the FNO, it seamlessly outputs the wavefield for the larger domain, which we refer to as ’direct prediction’, as depicted in Figure 10. Although the main components of the wavefield have been predicted by the FNO, their relative errors are still large. We calculate the relative L2 norm, which is the Euclidean distance from the prediction to the reference solutions divided by the Euclidean norm of the reference solutions, and the relative errors for direct predictions are 0.542 for the real part and 0.537 for the imaginary part of the wavefields. It is hard to obtain a high-accuracy prediction for a large unseen domain directly. Crucially, leveraging the reference frequency concept (Huang & Alkhalifah, 2022b) allows us to upscale the 4.78 Hz wavefield in the large domain (12.8$\times$12.8 km2) to an equivalent 9.56 Hz wavefield in the smaller domain (6.4$\times$6.4 km2). This provides us with physical support to predict high-frequency wavefields for a large-scale velocity model rather than a direct prediction of high-resolution wavefields. As illustrated in Figure 11, this scaled wavefield demonstrates enhanced accuracy compared to direct predictions. The relative errors calculated at the resolution of 256$\times$256 between the predictions of our method without interpolation and the reference results are 0.143 for the real part and 0.142 for the imaginary part of the wavefield. After bilinear interpolation of the prediction using our method, the errors at a resolution of 512$\times$512 are 0.171 for the real part and 0.170 for the imaginary part of the wavefield. Though the errors slightly increase, as shown in Figure 11, the predictions using our method are still close to the reference results.
@@
Figure 11: A generalization test on a realistic velocity model (the same velocity shown in Figure 10) using our method. Columns 1, 2, and 3 show the numerical solutions, the predictions of our method, and the errors.
4 Discussions
@@
The proposed method has shown good performance on the OpenFWI and realistic velocity models. It frees us at inference time from numerical simulations, in which the computational complexity is high. Taking the test on realistic models as an example, the numerical simulation for each velocity and frequency, executed in parallel on an Intel(R) Xeon(R) Gold 6230R CPU @2.10GHz, at a resolution of 139$\times$139 requires 1.27 seconds, and that at a resolution of 256$\times$256 and 512$\times$512 require 2.79 and 8.64 seconds, respectively. Using our method, the elapsed time on an NVIDIA A100 GPU (we do not fully utilize the GPU due to the relatively small input size and batch size) is 0.008 seconds for the resolution of 139$\times$139, 0.012 seconds for the resolution of 256$\times$256, and 0.031 seconds for the resolution of 512$\times$512. The main computationally intensive part of the proposed method is the training. In our experiments, it took 15.75 hours for training the neural operator for the ”CurveVelA velocity models, and 29.25 hours for the realisitc velocity models. Noteworthy, all of the training are performed offline. When trying to apply our method, indeed the neural operator-based wavefield solutions, on the downstream tasks such as reverse time migration or full waveform inversion, which requires massive simulations on different frequencies and source locations, the speed-up ratio of our method compared to conventional numerical methods will significantly increase, which will easily justify the expensive training. Taking multiple frequencies simulations on a resolution of 139$\times$139 as an example, the elapsed times for learned wavefield solutions and numerical finite-difference method with respect to the number of frequencies are shown in Figure 12. Since the learned wavefield solutions make it easy to parallel the prediction without additional modification (like parallel computing required for the numerical method to handle multiple frequencies simultaneously), instead, we just increase the batch size of the input tensors. It is obvious that within 16 simulations, the elapsed time for the learned wavefield solutions did not exceed 1 s, while the elapsed time for the numerical simulation without additional modification increased linearly. Due to the saturation of computational resources and memory bandwidth, the cost of neural operator-based wavefield solutions is slightly increased, causing a non-linear increase in the speed-up ratio. In summary, at inference, it is much faster than the conventional numerical solvers. Those speedups come from two main aspects: (1) we avoid repeated simulations; (2) we can easily utilize the latest hardware and software advances of accelerated computing (Azizzadenesheli et al., 2024) because we build the algorithm on machine learning. However, we recognize the challenges associated with the preparation of data-label pairs required for training the neural operator, particularly as the resolution of training samples increases or when dealing with complex 3D cases, e.g., 3D viscoelastic anisotropic modeling. In addition, when dealing with high-resolution 3D simulation, the computational cost of FNO will remain a challenge. This might require light 3D neural networks such as 3D Unet (Wandel et al., 2021) or LordNet3D (Huang et al., 2024), which are efficient for 3D simulation.
Figure 12: The efficiency comparison in terms of elapsed time for the learned wavefield solutions and the numerical finite-difference method on a resolution of 139$\times$139 with respect to different numbers of frequencies.
@@
Figure 13: The relative L2 norm errors for each class of models. Regarding the sample index from 0 to 120, each of the 24 consecutive numbers represents a velocity model, and each of the three consecutive numbers represents a frequency. Then, each velocity model includes 8 frequencies, increasing in sequence, and each frequency value corresponds to three different shots.
@@
Finally, to further highlight the effectiveness of the proposed method on different neural network architectures, we compare the different approaches such as such as conventional inputs with full wavefield outputs (marked as ”Conventional in full out”), conventional inputs with scattered wavefield outputs (marked as ”Conventional in residual out”), our method with full wavefield outputs, and our method with scattered wavefield outputs on different neural network architectures. We pick ResNet (He et al., 2016) and Unet (Wandel et al., 2020) for the study. The results are shown in Tables 2 and 3. The results demonstrate that our method is generally effective with other neural networks. Regarding the type of output, we found that a residual output, given by the scattered wavefield here, will slightly improve the accuracy of the wavefield prediction when using U-Net while performing almost the same in ResNet and FNO. However, we found that for conventional input, which includes sub-optimal binary masks for the source location and constant channel for the frequency information, the residual output instead performs worse. We do not further investigate it in our paper and leave it for future investigation. Overall, these results indicate that the conventional, sub-optimal input is highly sensitive to the output type, resulting in inconsistent accuracy performance. In contrast, our proposed informative input, which includes the background wavefield and velocity as input, provides an additional benefit by making the neural operator-based simulation more robust to variations in the output type compared to the conventional approach.
Table 2: The training cost and accuracy comparison between different approaches, using ResNet as backbone
@@
Table 3: The training cost and accuracy comparison between different approaches, using Unet as backbone
@@
5 Conclusions
We presented a novel implementation for predicting the wavefield using the Fourier Neural Operator (FNO) by incorporating the analytically derived background wavefield as input. As a result, the source location and frequency are embedded in the background wavefield, facilitating compact and effective feature engineering. This methodology enables efficient surrogate modeling across various velocities, frequencies, and source locations. Our tests on both the OpenFWI dataset and some realistic velocity models have demonstrated the high accuracy and effectiveness of the proposed method. Additionally, incorporating a reference frequency strategy significantly enhances the accuracy of our model in generalizing to a larger, unseen domain, marking a substantial advancement in the field of frequency-domain wavefield modeling.
Acknowledgements.